A sample of quarterly amounts invested in Ethan-Jaden Company by its employees is depicted in the table below;


1
Expert's answer
2021-07-11T17:15:03-0400



We have to find

(i) How many employees invest in the company

(ii) Estimate the mean amount invested


Solution:

(i) Total number of employees invest in the company = "4+8+24+44+20+10+2+6=118"


(ii) Mean amount invested "\bar X = \dfrac{\sum f_ix_i}{\sum f_i}"





Now, Mean "\bar X = A+ \dfrac{\sum fd}{n}\cdot h"

"= 52.5+\dfrac{(-100)}{118}\cdot 5\\\ \\=52.5+(-0.8475\cdot 5)\\=52.5-4.2373\\=48.2627"